Find the time required for an investment of 5000 dollars to grow to 7200 dollars at an interest rate of 7.5 percent per year, compounded quarterly.
Found 2 solutions by sravanik2005, lwsshak3:
Answer by sravanik2005(14)

A = P(1+ r/n)^nt
here P is principal amount=5000
r=interest rate as decimal=0.075
and n=number of times the interest is compounded per year=4
t= number of years=? and A=amount of money accumulated after t years, including interest=7200 dollars
substitute these values in that formula
7200=5000(1+0.075/4)^(4t)
7200/5000=1.01875^4t => 1.01875^4t=36/25
ln(1.01875)^4t=ln (36/25)
4t(ln 1.01875)=ln(36/25)
=>t= ln(36/25)/(4ln 1.01875)=4.91 that is approximately 5 years

Answer by lwsshak3(11628)

Find the time required for an investment of 5000 dollars to grow to 7200 dollars at an interest rate of 7.5 percent per year, compounded quarterly.
**
compound interest formula: A=P(1+r)^n, P=initial investment, r=interest rate/period, n=number of periods, A=amt after n-periods.
For given problem:
P=5000
r=.075/4=.01875
n=to find
A=7200
..
7200=5000(1+.01875)^n
7200/5000=(1.0175)^n
1.44=(1.0175)^n
take log of both sides
log(1.44)=n*log(1.0175)
n=log1.44/log(1.0175)
n≈21 quarters or 5.25 years